OpenStudy (purplemexican):

A card is drawn from a standard deck of cards. Determine whether the events are mutually exclusive or inclusive. Then, find the probability. P(a red card or a queen)

OpenStudy (moazzam07):

hey aren't mutually exclusive: a red card can very well be a queen P(red card or queen) = P(red) + P(queen) - P(red & queen) = (26+4-2)/52 = 28/52 = 7/13 <----------- QED

OpenStudy (moazzam07):

what are the choices?

OpenStudy (purplemexican):

inclusive, 15/26 mutually exclusive, 15/26 inclusive, 7/13 mutually exclusive, 7/13

OpenStudy (moazzam07):

i dont see the right answer choice

OpenStudy (purplemexican):

inclusive right?

OpenStudy (moazzam07):

Mutually Inclusive is the right answer